# Algorithmic idealism: what should you believe to experience next?

### v1

@@ IV Algorithmic Idealism in a nutshell
How do we respond to Restriction A? One option is certainly to double down on physicalism, and to insist on saying that “what should I expect to experience next?” is not usually a well-defined question. In this view, we should see this as a kind of pragmatic question that sometimes does not have an answer, and that will often have a vague answer, similarly as the question “Are viruses alive?”
However, here I suggest to explore a second option, which is to assume that (a specific version of) this first-person question does always have an objective answer, and to work out a theory that aims at supplying this answer at least in principle. The motivation to do so is two-fold: pragmatically, we will at some point be in desperate need of guidance in exotic situations such as the one of Section I, and a precise theory that offers such guidance and that is also in agreement with the standard physics predictions in mundane situations seems to be a good tool to aim for. Furthermore, as I have argued in Section II, Quantum Theory suggests that the question of “what should I expect to experience next?” is a more fundamental, natural and fruitful question to ask than “what is the case in the world”. We may see this as a hint from modern physics that we can hope to get closer to the truth by attempting to answer the former rather than the latter, and we can hope that trying to do so may indirectly tell us something interesting about the world, too.
@@
At first sight, this seems absurd: after all, will the agent not make another observation soon that will either confirm or disprove its concerns of being a Boltzmann brain, and hence retrospectively say that it was actually the copy on the planet, and not the random fluctuation out there? On second thought, however, the apparent contradiction gets resolved by two insights. First, Algorithmic Idealism is constructed to avoid the “standard methodology” described in Section III, and to allow agents to predict what they will see next without grounding this on the expected state of the world into which they are embedded. Hence, by construction, what they should believe cannot depend on whether they are “actually” surrounded by a planet-like environment or high-entropy radiation. Second, as will be explained in Section V, Algorithmic Idealism predicts that agents can often think of themselves as “effectively embedded”: what happens to them will often look to them to excellent approximation as if they were a part of some external world with algorithmically simple, probabilistic laws of nature. In the cosmology example, this results in the consequence that the agent will probably experience “planetary business as usual” next, regardless of the number of Boltzmann brains, even though the question of whether it actually is a Boltzmann brain or not is ill-defined. This will be discussed in more detail in Section V below.
@@
Postulate 2 (State Change). If you are in self state $x$ now, you will be in another self state $y$ next, and which one this will be manifests itself as a random experiment for you. Moreover, the likelihood of any $y$ next, given $x$ now, is what would be determined by some universal method of induction.
@@
Induction, broadly speaking, is a method to predict future observations by extrapolating regularities of past observations. What I mean by “universal” is a method that does not only work under the assumptions that we navigate within our own universe, but in a large class of conceivable environments. The possibility of inductive inference is one of the main assumptions of the scientific method. Here, we stipulate that induction should in principle always be possible, even in exotic situations, but in the following indirect way: there is a method of induction $M$ such that the likelihood of becoming $y$ next, given that one is $x$ now, is exactly the one that is predicted by $M$. In other words, a hypothetical third person that uses method $M$ could correctly (albeit not deterministically) predict any other agent’s future self state, if it were given a complete description of that agent’s current self state. Another way to phrase Postulate 2 would be to say that regularities that are present in your current self state will tend to persist in your private future.
@@
Note that this does not imply that the agent itself (which is in self state $x$ now) could use method $M$ to predict its own future experiences directly. After all, the agent will not in general have complete knowledge of $x$, and some self states will represent observers that are incapable of reasoning, or that are simply patterns which defy any interpretation as “observer” or “agent”. However, agents like us who are able to reason rationally may still use the resulting theory, Algorithmic Idealism, to predict something about their future observations. This can be done either by working out general predictions of Algorithmic Idealism that are independent of the specific form of the current self state $x$, or by analyzing the consequences of some (incomplete, but non-zero) knowledge of $x$.
My goal in all of this is to obtain a formal, rigorously defined version of Algorithmic Idealism, and to do so, one needs to specify Postulates 1 and 2 in complete mathematical detail. There are obviously many ways in which this can be done: for every choice of self states $\mathcal{S}$ and method of induction $M$, one obtains a different possible theory. Furthermore, even the mathematical definition of what “likelihood” means (a probability distribution, or something more general?) may vary between different formalizations. One preliminary elaboration of this (the “bit model”) has been worked out in [22], and an improved version is currently in preparation. Algorithmic idealism should therefore be seen as a kind of conceptual framework rather than a single theory.
@@
That this is a good method of prediction can be shown as follows [35, Corollary 5.2.1]. Suppose that the bits that you see are actually generated by some computable process $\mu$ (a “recursive measure” in the terminology of [35]), which you do not know. Then, with $\mu$-probability one,
That is, in the long run, algorithmic probability $\mathbf{M}$ will almost surely agree with the actual probabilities $\mu$ determined by the unknown computable process that generates the bits that you see.
@@
This extremely simple model is quite powerful: it admits the rigorous proof of several important predictions that I will describe in the following section. However, it has some undesirable features that motivate my current work on an improved version. In particular, it implies that a momentary state $x$ has always complete information about what the previous self states have been (namely, the prefixes of $x$). This is both unrealistic (thinking, say, of human observers) and unmotivated. It is simply an artefact of the way that AIT has been pursued so far.
@@
V Emergent objective reality, probabilistic changelings, and other predictions
@@
In this section, I will argue that Algorithmic Idealism can satisfy all these desiderata. I will do so via plausibility arguments, based on Postulates 1 and 2 of Section IV. In [22], I have given rigorous formulations and proofs of these claims for the bit model as described above. However, for the reasons mentioned above, the bit model has to be improved. My hope is that the colloquial argumentation of the present paper can be a useful guide for the proof of the claims in any mathematical formalization of Algorithmic Idealism. It certainly gives a valid conceptual summary of what is going on in the technical proofs for the bit model of [22].
@@
The main argument for why Algorithmic Idealism will be in agreement with our physical theories is that physics seems to predict that induction is possible. Consider an agent characterized by some self state $x$ which contains a lot of data. As we have said above, a self state need not correspond to anything that we typically interpret as an agent or observer, so let us consider planet Earth. Motivated by AIT, we may focus on versions of Algorithmic Idealism where the set of all possible self states is the finite binary strings, $\mathcal{S}=\{\varepsilon,0,1,00,01,\ldots\}$. To relate Earth with a binary string, let us fix an arbitrary method for encoding an approximate description of a momentary [^13] We may have to decide how to deal with relativity’s lack of absolute simultaneity and other ambiguities, but any decision for how to do so will considered to be part of the definition of the encoding procedure. configuration of Earth into bits. For example, we might coarsegrain the shell of the sea level plus or minus one kilometer into approximate cubes of sidelength $10^{-9}$ meters, and store a couple of bits for each cube that describe the approximate material composition of the corresponding cube, resulting in an enormous chunk of data.
@@
For the sake of the argument, imagine that we give this data to a third party that has access to astronomical computational power, so that it is able to implement a universal method $M$ of inductive inference. We do not tell this third party where the data is coming from, but we ask it to make a prediction for what this source of data will yield next. (Even if we think of continuous time evolution in our world, the result of this discrete encoding procedure will discretely change in a non-zero but very small fraction of a second in the future, and this change is what the third party is supposed to predict.) Earth contains a lot of records about the universe in which it is located: there are fossils in the underground implying biological evolution, astronomical photographs in libraries that convey some knowledge about what the world looks like external to Earth, and movies or other subtle properties of its structure that reveal many aspects of the physical laws in the relevant regime. It is not unreasonable to expect that a complete description of the relevant laws of physics and, say, the relevant aspects of the state of the universe at the Big Bang can be extracted from this data In Principle, which is capital I and P because it is unimaginably far removed from in practice. A very good method $M$ of inductive inference should In Principle discover all these regularities in the data, and predict whatever our universe’s laws and initial state would imply, including probabilities that we may think of as arising from a quantum description.
@@
If we have AIT in mind, then this method of induction could be something similar to Solomonoff induction (indeed, it is exactly this in the bit model of [22]). As we have seen in the previous section, Solomonoff induction gives us asymptotically correct probabilistic predictions for any probabilistic process that is computable. But the laws of physics as we know them are computable in some sense, which is described in more detail in physical versions of the Church-Turing thesis [36, 37, 38, 39, 40, 41], see also [22, p. 11, footnote 8]. And this seems to imply that finite-data processes such as the one described above (discretized Earth) are computable in the right technical sense to conclude that Solomonoff induction [^14] To apply Solomonoff induction in its standard formulation, we would have to set up the encoding such that the binary string can only grow over time by appending new bits. What I have in mind here is another method of induction that does without this assumption, which is currently work in progress. $M$ gives very similar predictions as our (best contemporary or even better future) physical theories.
@@
In the above, there was nothing special about Earth or about the specific choice of encoding procedure. We could instead think of anything else that contains a large amount of data about our world, for example a human brain, a hard disc, or the left half of the brain of some mammal, to some level of detail of description. In all those cases, Postulate 2 should yield very similar (and asymptotically identical) predictions for what happens to you if you are that thing [^15] To be more precise, it is not about being “that thing”, but about being the self state that is encoded into it. as physics does.
@@
Algorithmic Idealism seems to predict this sort of consistency with physics, but all the above reasoning says is that agents who appear to be embedded in a simple probabilistic universe are also likely to make observations in the future that confirm this. But why should this former premise be the case in the first place, if Algorithmic Idealism is true? After all, one of the main methodological principles of Algorithmic Idealism is to drop the assumption that an agent is fundamentally embedded into an external world. So how can we understand why agents like us are effectively embedded, i.e. that what happens to us appears as if we were embedded?
@@
One main insight shown for the bit model, but hopefully true much more generally, is the following somewhat technical result within AIT. Algorithmic probability measures such as $\mathbf{M}$ are not computable. However, they may often behave very similarly as a computable measure $\mu$, in particular if $\mu$ is algorithmically simple. But if this turns out to be the case, then the evolution of the agent’s self state will approximately be described by $\mu$, and there will by definition exist [^16] This is a statement of mathematical existence, not a claim of actual physical implementation, whatever the latter would mean. In particular, Algorithmic Idealism does not claim that there are actual machines that perform the computations which will naturally arise as external worlds for our experiences; these computations are in some sense extremely convenient (and hence meaningful) fictions. a computation, based on a short program (“simple laws of nature”), that generates $\mu$. To predict what happens to the agent, we may hence pretend that it is part of a computational process that generates $\mu$, and this process plays the role of an external world.
@@
Colloquially, we can think of $K(\mu)$ being small ($2^{-K(\mu)}$ being large) as saying that there is a simple probabilistic computational process that generates $\mu$. It can be modelled as running on some machine with a dedicated part (for Turing machines [^17] Algorithmic Idealism does not tell you what the computational model is; all computational models can simulate each other and are hence algorithmically equivalent, up to qualifications such as “being prefix-free” which have direct impact on the set of all possible algorithmic probability distributions that these machines generate. But whether you, say, describe an algorithm in terms of a Turing machine or of a cellular automaton is completely up to you. In particular, Algorithmic Idealism does not claim that an agent’s external world should resemble any of our mathematical models that we happen to have accidentally introduced and studied in computer science., the output tape) that contains the self state $x$ which evolves when the computation unfolds. As every computation, it starts in some simple initial state, with a short program of length $K(\mu)$ that tells the machine what to do. After this, it unfolds probabilistically, and it contains many other things in addition to the self state (for Turing machines, it would be other tapes, working memory etc.) such that the resulting evolution of the self state can be understood to arise from this in a “mechanical” manner.
But this is a very good description of what we mean by our external world $W$: a larger process into which we seem embedded, and which seems to have once been in a simple distinguished state (the Big Bang) and then evolved probabilistically according to simple laws. We can hypothesize about things in our external world (e.g., a brick falling towards my foot) and use this to construct mechanistic, causal explanations for the evolution of our self state (e.g., the sensation of pain and the observation that something red can be seen when I look down). The fact that our world is formally computable in some sense, and can hence colloquially be understood as a computation, is a well-known and widely discussed thesis that comes in different versions [50, 51].

### v2

@@ IV Algorithmic idealism in a nutshell
How do we respond to Restriction A? In the case of $N\geq 2$ agents, the nonexistence of a joint probability distribution for the observations of all agents can be interpreted [^15] It certainly does not have to be interpreted in this way (e.g., one might drop some background assumptions such as locality that underly the corresponding no-go theorems), but it is a well-motivated and natural possibility. as evidence that we should not always regard observations or events as absolute (see e.g. [25, 39, 40]), and algorithmic idealism as described below will agree with this. The case of Restriction A for a single agent, $N=1$, is however even more troublesome: it means that an agent can perform an experiment where they will learn some outcome, but there is nothing that can be said about what they should expect to observe. As explained further above, I regard this as absurd, and interpret it as a shortcoming of our given theories. If one takes this viewpoint, how does one react to it?
One option is certainly to double down on physicalism, and to insist on saying that “what should I expect to experience next?” is not usually a well-defined question. In this view, we should see this as a kind of pragmatic question that sometimes does not have an answer, and that will often have a vague answer, similarly as the question “Are viruses alive?”
However, here I suggest to explore a second option, which is to assume that (a specific version of) this first-person question does always have an objective answer, and to work out a theory that aims at supplying this answer at least in principle. The motivation to do so is two-fold: pragmatically, we will at some point be in desperate need of guidance in exotic situations such as the one of Section I, and a precise theory that offers such guidance and that is also in agreement with the standard physics predictions in mundane situations seems to be a good tool to aim for. Furthermore, as I have argued in Section II, Quantum Theory suggests that the question of “what should I expect to observe next?” is a more fundamental, natural and fruitful question to ask than “what is the case in the world”. We may see this as a hint from modern physics that we can hope to get closer to the truth by attempting to answer the former rather than the latter, and we can hope that trying to do so may indirectly tell us something interesting about the world, too.
@@
At first sight, this seems absurd: after all, will the agent not make another observation soon that will either confirm or disprove its concerns of being a Boltzmann brain, and hence retrospectively say that it was actually the copy on the planet, and not the random fluctuation out there? On second thought, however, the apparent contradiction gets resolved by two insights. First, algorithmic idealism is constructed to avoid the “standard methodology” described in Section III, and to allow agents to predict what they will see next without grounding this on the expected state of the world into which they are embedded. Hence, by construction, what they should believe cannot depend on whether they are “actually” surrounded by a planet-like environment or high-entropy radiation. Second, as will be explained in Section VI, algorithmic idealism predicts that agents can often think of themselves as “effectively embedded”: what happens to them will often look to them to excellent approximation as if they were a part of some external world with algorithmically simple, probabilistic laws of nature. In the cosmology example, this results in the consequence that the agent will probably experience “planetary business as usual” next, regardless of the number of Boltzmann brains, even though the question of whether it actually is a Boltzmann brain or not is ill-defined. This will be discussed in more detail in Section VI below.
@@
Postulate 2 (State Change). If you are in self state $x$ now, you will be in another self state $y$ next, and which one this will be manifests itself as a random experiment for you. The likelihood of some $y$ next, given $x$ now, has an objective value, which expresses what you should believe will happen to you next, if you knew your current self state. The mathematical structure that defines its value formalizes a universal method $M$ of induction.
@@
Induction, broadly speaking, is a method to predict future observations by extrapolating regularities of past observations. What I mean by “universal” is a method that does not only work under the assumptions that we navigate within our own universe, but in a large class of conceivable environments. The possibility of inductive inference is one of the main assumptions of the scientific method. Here, we stipulate that induction should in principle always be possible, even in exotic situations, but in the following indirect way: there is a method of induction $M$ such that the likelihood of becoming $y$ next, given that one is $x$ now, is exactly the one that is predicted by $M$. In other words, a hypothetical third person that uses method $M$ could correctly (albeit not deterministically) predict any other agent’s future self state, if it were given a complete description of that agent’s current self state. A consequence of Postulate 2 is that regularities that are present in your current self state will tend to persist in your private future, because methods of induction assign high probability to this being true.
@@
Probabilities obtained by some method of induction are naturally interpreted in a Bayesian way: many such methods begin with a certain choice of prior, and apply Bayesian updating upon obtaining new information (this is in particular true of Solomonoff Induction, used in the bit model below). However, here, the mathematical formulation of a method of induction is detached from its epistemic origins: rather than describing someone’s beliefs about an unknown future, the values that it defines mathematically are now interpreted as some sort of objective chance of turning from some given self state to another one. In more detail, they are supposed to be interpreted in a similar way as Berghofer’s DEJI interprets quantum probabilities: not as expressions of what somebody believes, but what they should believe if they were told the current self state [^18] This does not imply that the agent itself (which is in self state $x$ now) can use method $M$ to predict its own future experiences directly. After all, the agent will not in general have complete knowledge of $x$, and some self states will represent observers that are incapable of reasoning, or that are simply patterns which defy any interpretation as “observer” or “agent”. However, agents like us who are able to reason rationally may still use the resulting model of algorithmic idealism to predict something about their future observations. This can be done either by working out general predictions of algorithmic idealism that are independent of the specific form of the current self state $x$, or by analyzing the consequences of some (incomplete, but non-zero) knowledge of $x$..
Postulates 1 and 2 are intentionally formulated in such a way that they admit of several possible rigorous mathematical formulations, or models. For every choice of self states $\mathcal{S}$ and method of induction $M$, one obtains a different model. One simple, preliminary model, the bit model, has been introduced in [2]. It will be described in Section V below. Different mathematical models (and the associated methods of induction) may rely on different mathematical definitions of likelihood, e.g. on a single conditional probability distribution, on a collection of several such distributions (which is the case for the bit model below), on a notion of imprecise probability [29], or on plausibility measures [30, 31]. The precise conceptual interpretation of these “values” of likelihood, as currently sketched in Postulate 2, may have to be adapted on a case-by-case basis. Algorithmic idealism should therefore be seen as a kind of conceptual framework rather than a single theory, because it admits many different mathematical realizations.
@@ V Algorithmic information theory and the bit model
That this is a good method of prediction can be shown as follows [45, Corollary 5.2.1]. Suppose that the bits that you see are actually generated by some computable process $\mu$ (a “recursive measure” in the terminology of [45]), which you do not know. Then, with $\mu$-probability one,
$\lim_{n\to\infty}\left|\mathbf{M}(b|x_{1}^{n})-\mu(b|x_{1}^{n})\right|=0\qquad(b\in\{0,1\}).$ (2)
That is, in the long run, algorithmic probability $\mathbf{M}$ will almost surely agree with the actual probabilities $\mu$ determined by the unknown computable process that generates the bits that you see. For example, suppose that the process always outputs deterministically the bit $1$, i.e. $\mu(1^{n})=1$ for all $n\in\mathbb{N}$, where $1^{n}=\underbrace{11\ldots 1}_{n}$. Then $\mathbf{M}(0|1^{n})=2^{-K(n)+\mathcal{O}(1)}$, which tends to zero and is of the order $1/n$ for most $n$. In this sense, after seeing $n$ ones, Solomonoff induction predicts that you will probably not see a zero next, as expected from a method of induction.
@@
This extremely simple model is quite powerful: it admits the rigorous proof of several significant predictions that I will describe in the following section. However, it has some undesirable features that motivate my current work on an improved version. In particular, it implies that a momentary state $x$ has always complete information about what the previous self states have been (namely, the prefixes of $x$). This is both unrealistic (thinking, say, of human observers) and unmotivated. It is simply an artefact of the way that AIT has been pursued so far.
@@
VI Predictions of algorithmic idealism
@@
In this section, I will argue that algorithmic idealism can satisfy all these desiderata. I will do so via plausibility arguments, based on Postulates 1 and 2 of Section IV. In [2], I have given rigorous formulations and proofs of these claims for the bit model as described above. However, for the reasons mentioned above, the bit model has to be improved. My hope is that the colloquial argumentation of the present paper can be a useful guide for the proof of the claims in other mathematical formalizations of algorithmic idealism. It certainly gives a valid conceptual summary of what is going on in the technical proofs for the bit model of [2].
VI.1 Consistency with physical predictions in standard scenarios
@@
The main argument for why algorithmic idealism will be in agreement with our physical theories is that physics seems to predict that induction is possible. Consider an agent characterized by some self state $x$ which contains a lot of data. As we have said above, a self state need not correspond to anything that we typically interpret as an agent or observer, so let us consider planet Earth. Motivated by AIT, we may focus on versions of algorithmic idealism where the set of all possible self states is the finite binary strings, $\mathcal{S}=\{\varepsilon,0,1,00,01,\ldots\}$. To relate Earth with a binary string, let us fix an arbitrary method for encoding an approximate description of a momentary [^21] We may have to decide how to deal with relativity’s lack of absolute simultaneity and other ambiguities, but any decision for how to do so will considered to be part of the definition of the encoding procedure. Mathematically, the procedure must be computable, see the notion of a computational ontological model of the technical paper [2]. configuration of Earth into bits. For example, we might coarsegrain the shell of the sea level plus or minus one kilometer into approximate cubes of sidelength $10^{-9}$ meters, and store a couple of bits for each cube that describe the approximate material composition of the corresponding cube, resulting in an enormous chunk of data.
@@
For the sake of the argument, imagine that we give this data to a team of brillant scientists that has access to astronomical computational power. We ask the team to make a prediction for what this source of data will yield next. (Even if we think of continuous time evolution in our world, the result of this discrete encoding procedure will discretely change in a non-zero but very small fraction of a second in the future, and this change is what the third party is supposed to predict.) Earth contains a lot of records about the universe in which it is located: there are fossils in the underground implying biological evolution, astronomical photographs in libraries that convey some knowledge about what the world looks like external to Earth, and movies or other subtle properties of its structure that reveal many aspects of the physical laws in the relevant regime. It is not unreasonable to expect that a complete description of the relevant laws of physics and, say, the relevant aspects of the state of the universe at the Big Bang can be extracted from this data in principle – after all, this is what we humans seem to accomplish successfully via science. Hence, the team of scientists should be able to accomplish their task successfully.
What this thought experiment then shows is the first step in the following three-step argument:
  1. 1.
    We have heuristically argued, given the success of science, that good methods of induction – if they were supplied with a description of a momentary state of Earth as just described – should predict the same probabilities of future states of Earth as our best scientific theories. Hence if science makes correct predictions, good methods of induction will do so as well.
  2. 2.
    As a next step, we lift this from a pragmatic, heuristic, methodological intuition to a precise mathematical claim: if we are given a mathematically well-defined method of universal induction $M$ (such as Solomonoff induction defined above), then $M$ should also make correct predictions about the probabilities of future states of Earth. This is much less intuitive, since the input to $M$ is simply a bare self state; in the bit model, it is a long string of binary data. It may seem surprising (and for philosophically-minded readers highly dubious) that it is enough to have this data without any built-in semantics, such as a reference to spatiotemporal structure etc., and still to obtain correct predictions from it. But this is guaranteed by mathematical theorems (essentially eq. (2)) if the induced stochastic process on the self states is computable.
@@
    Indeed, the laws of physics as we know them are computable in some sense, which is described in more detail in physical versions of the Church-Turing thesis [46, 47, 48, 49, 50, 51], see also [2, p. 11, footnote 8]. And this seems to imply that finite-data processes such as the one described above (discretized Earth) are computable in the right technical sense to conclude that Solomonoff induction [^22] To apply the bit model, and hence Solomonoff induction in its standard formulation, we would have to set up the encoding such that the binary string can only grow over time by appending new bits. Constructing a model that does not suffer from this restriction is the subject of ongoing work. $M$ makes asymptotically correct predictions.
  3. 3.
    Steps 1 and 2 have only argued about general properties of universal induction, formulated via well-known results of algorithmic information theory. Algorithmic idealism becomes only relevant now: it says that if you are Earth’s current self state [^23] As discussed further above, self states are not in general associated to intuitive instances of “agents” or “observers”., then what you will become next is determined by $M$ (in the bit model, algorithmic probability). In particular, it is not determined by the laws of any world in which you might imagine being embedded (actually, you are unembedded). Instead, $M$ is the “law of nature”, and its action depends on the bare self state and not on anything external to it.
    What steps 1 and 2 now show is that algorithmic idealism’s claim of what happens to you if you are Earth (in more detail, if you are its data defined by the specific choice of read-out procedure) is consistent with what happens to Earth according to our usual physical picture of the world.
@@
In the above, there was nothing special about Earth or about the specific choice of encoding procedure. We could instead think of anything else that contains a large amount of data about our world, for example a human brain, a hard disc, or the left half of the brain of some mammal, to some level of detail of description. In all those cases, Postulate 2 should yield very similar (and asymptotically identical) claims for what happens to you if you are that thing [^24] To be more precise, it is not about being “that thing”, but about being the self state that is encoded into it. as physics does.
In the bit model, we can be more specific what the requirements are for this consistency result to hold. Consider some bit-string-valued random variable $x$ in the world (see the technical paper [2, Example A.1] for how this is consistent with quantum theory), changing over time in a way that its length does not decrease. For example, think of a video camera that is continuously filming Times Square and recording the pictures on a hard disc, where the video is encoded into bits. Or think of the “state of the Earth” example above, or a discretization (and arbitrary computable method of digitization) of your brain to some large accuracy over time, keeping all past information. Our physical theories tell us that the physical world generates a probability measure $\mu$ on the possible histories of this bit string (for example, pictures of the Times Square showing all pedestrians suddenly floating up towards the sky has extremely small probability). If a suitable version of a physical Church-Turing thesis is true, then $\mu$ will be a computable measure. (For this, not only the process defining the relevant aspects of the world, but also the definition of the random variable needs to be computable.) Then, Eq. (2) will be true: predictions via algorithmic idealism will in the long run agree with the predictions via physics.
Recall that Postulate 2 is not a description of what we (or the agent) believes will happen next, but what they should believe. That is, it is understood, in its rigorous formulation within a specific choice of mathematical model (such as the bit model), as a statistical law of nature. In particular, it is not means as a supplement of the given laws of physics, but as a replacement [^25] It would certainly be a very bad idea to try and use algorithmic idealism, or algorithmic probability, directly in the regime of physics. It will never be a good replacement in practice, it is only a very abstract theoretical (but important) fact that it could be in principle.. Therefore, it is crucial to have a theorem that guarantees its consistency with the standard physical laws in the regime where they apply. This is essentially the regime of empirical science, i.e. of processes where a large amount of data is collected over time.
VI.2 The external world as an emergent approximate phenomenon
@@
Algorithmic idealism predicts this sort of consistency with physics, but all the above reasoning says is that agents who appear to be embedded in a simple probabilistic universe are also likely to make observations in the future that confirm this. But why should this former premise be the case in the first place, if algorithmic idealism is true? After all, one of the main methodological principles of algorithmic idealism is to drop the assumption that an agent is fundamentally embedded into an external world. So how can we understand why agents like us are effectively embedded, i.e. that what happens to us appears as if we were embedded?
@@
One main insight shown for the bit model, but hopefully true much more generally, is the following somewhat technical result within AIT. Algorithmic probability measures such as $\mathbf{M}$ are not computable. However, they may often behave very similarly as a computable measure $\mu$, in particular if $\mu$ is algorithmically simple. This may in particular be the case if, by chance, the previously observed bits $x_{1}^{n}=x_{1}x_{2}\ldots x_{n}$ happen to correspond to a typical result of the probability measure $\mu$; intuitively, in this case, the following bits will also likely be distributed in accordance with $\mu$. Intuitively, the statistical regularities present in $x_{1}^{n}$ will tend to be preserved because this is what induction deems likely to be the case. This leads to a general tendency of “stabilizing” a very regular statistical behavior: one that is according to a simple computable measure $\mu$. If this turns out to be the case, then the evolution of the agent’s self state will approximately be described by $\mu$, and there will by definition exist [^26] This is a statement of mathematical existence, not a claim of actual physical implementation, whatever the latter would mean. In particular, algorithmic idealism does not claim that there are actual machines that perform the computations which will naturally arise as external worlds for our experiences; these computations are in some sense extremely convenient (and hence meaningful) fictions. a computation, based on a short program that generates $\mu$. To predict what happens to the agent, we may hence pretend that the agent is part of a computational process that generates its self state according to the computable distribution $\mu$, and this process plays the role of an external world.
@@
That is, for every computable measure $\mu$, there is a positive probability that algorithmic probability (and hence the actual chances of what happens to the agent), conditioned on the previous observations, will converge to $\mu$. Not all measures are equally likely, though. Colloquially, we can think of $K(\mu)$ being small (probability $2^{-K(\mu)}$ being large) as saying that there is a probabilistic computational process that generates $\mu$ which has a short program, i.e. is simple in this sense. It can be modelled as running on some machine with a dedicated part (for Turing machines [^27] Algorithmic idealism does not tell you what the computational model is; all computational models can simulate each other and are hence algorithmically equivalent, up to qualifications such as “being prefix-free” which have direct impact on the set of all possible algorithmic probability distributions that these machines generate. But whether you, say, describe an algorithm in terms of a Turing machine or of a cellular automaton is completely up to you. In particular, algorithmic idealism does not claim that an agent’s external world should resemble any of our mathematical models that we happen to have accidentally introduced and studied in computer science., the output tape) that contains the self state $x$ which evolves when the computation unfolds. As every computation, it starts in some simple initial state, with a short program of length $K(\mu)$ that tells the machine what to do. After this, it unfolds probabilistically, and it contains many other things in addition to the self state (for Turing machines, it would be other tapes, working memory etc.) such that the resulting evolution of the self state can be understood to arise from this in a “mechanical” manner.
But this is a very good description of what we mean by our external world $W$: a larger process into which we seem embedded, and which seems to have once been in a simple distinguished state (the Big Bang) and then evolved probabilistically according to simple laws. We can reason about things in our external world (e.g., a brick falling towards my foot) and use this to construct mechanistic, causal explanations for the evolution of our self state (e.g., the sensation of pain and the observation that something red can be seen when I look down). The fact that our world is formally computable in some sense, and can hence colloquially be understood as a computation, is a well-known and widely discussed thesis that comes in different versions [61, 62].
@@
VI.3 More than one observer: emergent objective $\ldots$